# Improved GUI Grounding via Iterative Narrowing

### v4

@@ Abstract
Graphical User Interface (GUI) grounding plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents. While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal. Recent studies have focused on fine-tuning these models specifically for one-shot GUI grounding, yielding significant improvements over baseline performance. We introduce a visual prompting framework that employs an iterative narrowing mechanism to further improve the performance of both general and fine-tuned models in GUI grounding. For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
@@ 1 Introduction
Figure 1: A visual demonstration of our method. On each non-final iteration, the model prediction (red cross) determines the next region (red box) to iterate.
@@ 2.1 Grounding Models
Cheng et al. [4] trained SeeClick, a GUI-specific grounding model and introduced the ScreenSpot benchmark, showing that grounding pre-training enhances task performance. Gou et al. [5] expanded on this by training a pixel-level model on a dataset of 10 million GUI elements, achieving up to 20% accuracy improvements and demonstrating effective human-like visual grounding. Wu et al. [9] introduced a cross-platform dataset of 13 million GUI elements and trained two models: OS-Atlas-Base-4B and OS-Atlas-Base-7B, improving performance across diverse benchmarks.
@@ 2.2 GUI Agents
Xie et al. [10] introduced the OSWorld benchmark, designed to evaluate an agent’s ability to complete tasks within GUI environments. One of the key categories tested is the screenshot-only scenario, where the model is provided solely with an image of the screen, without any auxiliary information. This scenario emphasizes the importance of strong grounding capabilities, as a common failure case observed by the authors stems from inadequate grounding performance.
Wu et al. [9] expanded on this challenge by integrating their pre-trained grounding model, OS-Atlas-Base-7B, with GPT-4o, demonstrating significant improvements over GPT-4o alone on the OSWorld benchmark. Their results further validate the crucial role of effective grounding in improving task completion in GUI environments.
Hong et al. [6] also addressed GUI grounding by training and evaluating CogAgent on the screenshot-only category of OSWorld. CogAgent was developed to function both as a grounding model as well as a standalone visual agent.
2.3 Iterative Visual Reasoning Beyond Convolutions
Chen et al. [1] proposed an iterative visual reasoning framework that extends beyond traditional CNNs by incorporating a local module with spatial memory and a global graph-reasoning module for spatial and semantic interactions. The global module builds a graph connecting knowledge, image regions, and class assignments, enabling iterative updates through cross-feeding between the modules to refine predictions.
2.4 Iterative Localization in Proposal-Free Referring Expression Grounding
Sun et al. [7] tackled proposal-free referring expression grounding introducing a similar mechanism called iterative shrinking, which is guided by a reinforcement learning agent to localize target objects based on query sentences.
This iterative approach parallels our method, though Sun et al. uses reinforcement learning, while our method relies on VLM predictions to refine the target.
@@ 3 Methodology
The cropping window dimensions are adjusted based on the image orientation. For landscape-oriented images, the width and height are halved at each iteration. In the case of portrait-oriented images, the width is scaled less aggressively by a factor of $1/1.2$, while the height is halved.
@@ 3.1 Pseudocode
      Predict normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$ relative to $\mathcal{C}_{k-1}$, from $\mathcal{Q}$
@@
      • $w_{k}\leftarrow\frac{w_{k-1}}{2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
@@
      • $w_{k}\leftarrow\frac{w_{k-1}}{1.2},\quad h_{k}\leftarrow\frac{h_{k-1}}{2}$
@@ 4.1 Evaluation Results
Models Baseline Ours
InternVL-2-4B 4.32 6.53
@@
Table 1: Overall average accuracy comparing baseline against our method on the ScreenSpot benchmark.
Models Text Icon/Widget
Baseline Ours Baseline Ours
InternVL-2-4B 9.16 14.65 1.64 2.18
@@
Models Text Icon/Widget
Baseline Ours Baseline Ours
InternVL-2-4B 4.64 8.76 4.29 2.14
@@
Models Text Icon/Widget
Baseline Ours Baseline Ours
InternVL-2-4B 0.87 5.22 0.10 2.91
@@
For the entire evaluation, we used $n=3$ iterations. While this choice provided promising results, we did not conduct in-depth investigations to determine an optimal value for $n$ or how aggressive each crop should be shrunken. Future work may explore the impact of varying $n$ and the extent of each cropping step on performance and inference time.
@@ 4.2 Weaknesses & Observations
A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues. GUI-specific models like OS-Atlas-Base-7B [9] can accurately identify such targets in the baseline, while our method often fails due to limited context (see Figure 2).
@@
Our method improves precision but at the cost of underperforming in context-dependent scenarios. This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B [3, 2], Qwen2-VL-7B [8]) versus OS-Atlas-Base-7B [9], which already has strong precision abilities from training. Addressing this limitation by maintaining context throughout the iterations could further enhance performance for all models.

### v5

@@ Abstract
Graphical User Interface (GUI) grounding plays a crucial role in enhancing the capabilities of Vision-Language Model (VLM) agents. While general VLMs, such as GPT-4V, demonstrate strong performance across various tasks, their proficiency in GUI grounding remains suboptimal. Recent studies have focused on fine-tuning these models specifically for zero-shot GUI grounding, yielding significant improvements over baseline performance. We introduce a visual prompting framework that employs an iterative narrowing mechanism to further improve the performance of both base and finetuned models in GUI grounding. For evaluation, we tested our method on a comprehensive benchmark comprising various UI platforms and provided the code to reproduce our results in this GitHub repository.
@@ 1 Introduction
Figure 1: A visual demonstration of our method. On each iteration, the model prediction (red cross) determines the next region (red box) to iterate.
@@ 2.1 Grounding Models
Cheng et al. [4] trained SeeClick, a finetuned grounding model, and introduced the ScreenSpot benchmark, demonstrating the benefits of grounding pre-training. Gou et al. [5] achieved up to 20% accuracy gains with a pixel-level model trained on 10 million GUI elements. Wu et al. [10] developed OS-Atlas models using a 13-million GUI element dataset, improving performance across benchmarks. Lin et al. [7] proposed ShowUI, a lightweight vision-language-action model optimized for GUI tasks using innovative token selection and streaming methods.
@@ 2.2 GUI Agents
Xie et al. [11] introduced the OSWorld benchmark to evaluate GUI task agents, emphasizing grounding in screenshot-only scenarios. Wu et al. [10] enhanced task performance by integrating OS-Atlas-Base-7B with GPT-4o. Hong et al. [6] addressed GUI grounding with CogAgent, designed as both a grounding model and visual agent.
2.3 Iterative Reasoning and Localization
Chen et al. [1] introduced a visual reasoning framework combining spatial memory and graph reasoning modules to iteratively refine predictions through spatial and semantic interactions. Similarly, Sun et al. [8] proposed an iterative shrinking approach for referring expression grounding using reinforcement learning, which parallels our method’s reliance on VLM predictions for target refinement.
@@ 3 Methodology
The cropping window dimensions are adjusted dynamically based on the image orientation. For landscape-oriented images, the width and height are scaled by factors $\alpha_{\text{landscape}}$ and $\beta_{\text{landscape}}$, respectively, where both scale factors are less than one. Similarly, for portrait-oriented images, the width is scaled by $\alpha_{\text{portrait}}$, which is less aggressive than the landscape scaling, while the height is scaled by $\beta_{\text{portrait}}$.
@@ 3.1 Pseudocode
      Using VLM: Predict normalized coordinates $(x_{k},y_{k})\in[0,1]\times[0,1]$ relative to $\mathcal{C}_{k-1}$, from $\mathcal{Q}$
@@
      • $w_{k}\leftarrow\alpha_{\text{landscape}}\cdot w_{k-1},\quad h_{k}\leftarrow\beta_{\text{landscape}}\cdot h_{k-1}$
@@
      • $w_{k}\leftarrow\alpha_{\text{portrait}}\cdot w_{k-1},\quad h_{k}\leftarrow\beta_{\text{portrait}}\cdot h_{k-1}$
@@ 4 Experiments
4.1 Setup
For the entire evaluation, we used $n=3$ iterations. While this choice provided promising results, we did not conduct in-depth investigations to determine an optimal value for $n$ or how aggressive each crop should be shrunken. The cropping aggressiveness was controlled using scaling factors:
  • For landscape-oriented images:
    • Width scaling factor: $\alpha_{\text{landscape}}=0.5$
    • Height scaling factor: $\beta_{\text{landscape}}=0.5$
  • For portrait-oriented images:
    • Width scaling factor: $\alpha_{\text{portrait}}=\frac{1}{1.2}\approx 0.833$
    • Height scaling factor: $\beta_{\text{portrait}}=0.5$
Future work may explore the impact of varying $n$ and the values of $\alpha_{\text{landscape}}$, $\beta_{\text{landscape}}$, $\alpha_{\text{portrait}}$, and $\beta_{\text{portrait}}$ on performance and inference time.
@@ 4.2 Evaluation Results
Model Type Models Baseline IN
Base InternVL-2-4B 4.32 6.53
@@
Table 1: Overall average accuracy (%) comparing baseline against our method (IN) on the ScreenSpot benchmark.
Model Type Models Text Icon/Widget
Baseline IN Baseline IN
Base InternVL-2-4B 9.16 14.65 1.64 2.18
@@
Model Type Models Text Icon/Widget
Baseline IN Baseline IN
Base InternVL-2-4B 4.64 8.76 4.29 2.14
@@
Model Type Models Text Icon/Widget
Baseline IN Baseline IN
Base InternVL-2-4B 0.87 5.22 0.10 2.91
@@ 4.3 Weaknesses & Observations
A key weakness of our method is the loss of contextual information as iterations progress, posing challenges for identifying target elements that rely on spatially distant cues (see Figure 2).
@@
Our method improves precision but at the cost of underperforming in context-dependent scenarios. This limitation likely explains the greater gains observed in generalist VLMs (e.g., InternVL-2-4B [3, 2], Qwen2-VL-7B [9]) versus OS-Atlas [10] and ShowUI [7], which already has strong precision abilities from training. Addressing this limitation by maintaining context throughout the iterations could further enhance performance for all models.